# Understanding Pre-training and Fine-tuning from Loss Landscape Perspective

### v1

@@ D.2 Loss Landscapes of Large Models
Larger models are more robust to fine-tuning and jailbreaking. As analyzed in Sec. 4.1, the larger the basin size, the greater the robustness against fine-tuning and jailbreaking attacks. Therefore, we conclude that larger models are more robust to fine-tuning, less likely to compromise capabilities, and more resistant to jailbreaking attacks. This may be one of the benefits of scaling up models.
D.3 Hypothesis Testing
To determine the basin size for a given model, we need to test Definition 4. However, this requires calculating an expectation $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$ over a high-dimensional space $d$. Following [18, 60, 14], we use hypothesis testing to control the type-I error in Monte Carlo sampling for estimating this expectation. Specifically, we apply the Clopper-Pearson bound to obtain a lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$ with a type-I error of 0.01.
D.4 About Norm Constraints in Landscape Visualization
@@
Appendix E Norm of Common Fine-tuning
@@
We also test the $\ell_{2}$ norm of common post-training techniques to determine whether they are located within basins or guaranteed regions. As shown in Table 1, we draw the following conclusions:
@@
Appendix F Comparison of Eq. 6 and Eq. 7
@@
Beyond the fact that Eq. 7 is strictly stronger than Eq. 6 (i.e., the Lipschitz constant of Eq. 7 at each point is strictly less than $\frac{1}{\sqrt{2\pi}\sigma}$), there is also a significant difference. Eq. 7 indicates that the higher the clean accuracy $p_{A}$, the smaller the Lipschitz constant, and thus the greater the robustness; when $p_{A}=1$, the Lipschitz constant becomes zero, making forgetting extremely unlikely. Conversely, the lower the clean accuracy $p_{A}$, the larger the Lipschitz constant, and the more likely performance degradation occurs. This eliminates the forgetting-learning tradeoff even along the worst-case direction, i.e., as long as the performance on old tasks exceeds that on new tasks, the smoothness constraints make it more likely to learn new tasks than to forget old tasks, and the performance degradation on old tasks is always less than the maximum performance gain on new tasks.
Appendix G Limitations and Future Work
Preference for Pre-training Over Fine-tuning. The primary limitation of this work is our inability to apply the Gaussian-augmented optimizer during pre-training. When using the Gaussian-augmented optimizer only during fine-tuning, although the basin of basic capabilities can be widened, the points within the basin perform similarly to a randomly initialized model trained directly on the fine-tuning dataset without pre-training. Consequently, its other capabilities lag behind those of the original model. We hypothesize that this could be addressed by applying the Gaussian-augmented optimizer during pre-training rather than fine-tuning. Recent studies suggest that the pre-training phase is when the model acquires its capabilities, while fine-tuning merely activates these capabilities [50]. Using the Gaussian-augmented optimizer during pre-training may identify basins where most points already possess these capabilities.
Preference for Larger Models Over Smaller Models. As analyzed in Sec. D.2, larger models exhibit greater expressive power and more readily form larger basins. The primary obstacle preventing us from applying the Gaussian-augmented optimizer to both pre-train and larger models is their high memory usage. For a 7B model, a single 80GB GPU is insufficient to train with a batch size of 1, requiring the use of current training frameworks, such as DeepSpeed or Megatron, with optimizer offload modes. However, these frameworks support only a limited set of optimizers, such as Adam or SGD, and do not accommodate custom optimizers.

### v2

@@ D.3 Loss Landscapes of Large Models
Larger models are more robust to fine-tuning and jailbreaking. As analyzed in Sec.˜4.1, the larger the basin size, the greater the robustness against fine-tuning and jailbreaking attacks. Therefore, we conclude that larger models are more robust to fine-tuning, less likely to compromise capabilities, and more resistant to jailbreaking attacks. This may be one of the benefits of scaling up the model size.
Algorithm 2 Normalization of Benchmark Values ($\mathcal{T}$)
0: Benchmark values $\{J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$
1: Compute minimum value: $\text{min\_val}\leftarrow\min_{\alpha}J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
2: Compute maximum value: $\text{max\_val}\leftarrow\max_{\alpha}J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$
3: Compute range: $\text{range\_val}\leftarrow\text{max\_val}-\text{min\_val}$
4: Initialize normalized values: $\text{normalized}\leftarrow[]$
5:for each $\text{value}\in\{J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})\}_{\alpha\in A}$ do
6:  Scale to $[0,1]$: $\text{scaled}\leftarrow(\text{value}-\text{min\_val})/\text{range\_val}$
7:  Invert value: $\text{inverted}\leftarrow 1-\text{scaled}$
8:  Append to output: $\text{normalized}\leftarrow\text{normalized}+[\text{inverted}]$
9:end for
10:return normalized
D.4 Hypothesis Testing with Clopper-Pearson Bound
Soft Definition of Basins. To verify the soft definition of basins (Definition 3), we aim to confirm whether $J_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\epsilon$. We achieve this using the Clopper-Pearson bound by computing a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ with a specified type-I error $\gamma$. Specifically, the Clopper-Pearson bound ensures that $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\in[p_{\text{lower}},p_{\text{upper}}]$ with probability at least $1-\gamma$. If $J_{f,\mathcal{D}}(\bm{\theta})-p_{\text{lower}}\leq\epsilon$, we can assert, with type-I error $\gamma$, that the soft basin condition holds.
Strict Definition of Basins. For the strict definition of basins (Sec.˜3.2), our goal is to estimate the proportion of directions where the normalized benchmark value is exactly 0, i.e., to compute a lower bound for $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. The Clopper-Pearson bound provides a confidence interval $p_{\text{lower}},p_{\text{upper}}]$ with type-I error $\gamma$, allowing us to assert that at least $p_{\text{lower}}\times 100\%$ of directions form a strict basin with probability $1-\gamma$.
Introduction to the Clopper-Pearson Bound. The Clopper-Pearson bound constructs exact confidence intervals for the success probability $p$ of a binomial distribution, ensuring strict control over the type-I error rate $\gamma$, i.e., the probability of rejecting a true null hypothesis. Unlike approximate methods (e.g., normal approximation), it leverages the binomial likelihood and the Beta distribution to compute precise bounds, making it ideal for estimating basin sizes in large language models where accurate error control is critical.
Formulation and Application. Let $X$ denote a binomial random variable representing the number of successes in $n$ independent Bernoulli trials, where a success indicates that the perturbed model $f_{\bm{\theta}+\bm{\epsilon}}$ retains performance comparable to the original model $f_{\bm{\theta}}$. For the strict basin (Sec.˜3.2), a success occurs when $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0$. For the soft basin (Eq.˜3), since $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ is also an expectation of the Bernoulli variable (i.e., 0-1 loss on each instance) over $\mathcal{D}$, a success occurs when the answer of a sampled instance is judged as "correct" or "safe". Given a confidence level $1-\gamma$ (we use $\gamma=0.01$), the Clopper-Pearson confidence interval $p_{\text{lower}},p_{\text{upper}}]$ for $p$ is defined by:
$P(X\geq x\mid p=p_{\text{lower}})$ $=\frac{\gamma}{2},$ (9)
$P(X\leq x\mid p=p_{\text{upper}})$ $=\frac{\gamma}{2},$ (10)
where $x$ is the observed number of successes in $n$ trials. These bounds are computed using the inverse of the regularized incomplete Beta function:
$p_{\text{lower}}$ $=I^{-1}_{x,n-x+1}\left(\frac{\gamma}{2}\right),$ (11)
$p_{\text{upper}}$ $=I^{-1}_{x+1,n-x}\left(1-\frac{\gamma}{2}\right),$ (12)
where $I_{a,b}(z)$ is the cumulative distribution function of the Beta distribution $\text{Beta}(a,b)$. This ensures that our estimation of basin size maintains a type-I error rate below 0.01, providing a robust foundation for theoretical and experimental analyses.
Experimental Settings. Following Cohen et al. (2019); Salman et al. (2019); Chen et al. (2025a), we use Monte Carlo sampling with the Clopper-Pearson bound to estimate basin sizes. We set a type-I error of $\gamma=0.01$ and a sample size of $n=100,000$ to ensure precise confidence intervals. For the soft basin, we evaluate $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and $\sigma=0.01$, and obtain the statistical lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. For the strict basin, we sample directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ and compute $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ with $\alpha\in[0,3\times 10^{-3}]$, as in Table˜1. These settings enable reliable validation of the most-case basin properties, as visualized in Fig.˜1.
(a) NLL Landscape
(b) Hypothesis Testing
Figure 8: Loss landscape and hypothesis testing results using a likelihood-based benchmark. (a) The loss landscape of Qwen2.5-7B for a likelihood-based benchmark is smooth and continuous. (b) The lower bound $p_{\text{lower}}$ from the Clopper-Pearson bound for the soft basin definition (i.e., the lower bound for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$) and the strict basin definition (i.e., the lower bound for proportion of directions achieving the exact zero loss).
Experimental Results. The statistical lower bounds obtained for both the strict and soft definitions of basins are shown in Fig.˜8(b). As demonstrated, the lower bound for the soft basin definition significantly exceeds that of the strict basin definition, aligning with our observation that when $\epsilon\to 0$ in Definition 3, the soft definition becomes the strict definition of basins, assuming the expected loss does not decrease after adding noise. We present the respective conclusions for the two definitions as follows:
Results for Strict Definition of Basins. In the strict definition of basins, the lower bound $p_{\text{lower}}$ represents the proportion of directions forming strict basins. As shown in Fig.˜8(b), for $\sigma=0.01$, we can assert that more than 90% of directions form strict basins, whereas for $\sigma=0.02$, we can assert that more than 65% of directions form strict basins. As $\sigma$ increases, the proportion of directions that form strict basins decreases.
Results for Soft Definition of Basins. In the soft definition of basins, the lower bound $p_{\text{lower}}$ provides a statistical guarantee for $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$. As demonstrated, larger $\sigma$ values lead to worse model performance under noise perturbations but reduce the Lipschitz constant of the smoothed model. This reflects a classical robustness-accuracy tradeoff in randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Chen et al., 2024). Choosing a larger $\sigma$ enhances model robustness but may compromise accuracy, potentially resulting in either improved or degraded performance. In Sec.˜4, we select $\sigma=0.03$, for which $p_{\text{lower}}$ remains above 0.9.
D.5 Pretraining Using Go Optimizer
Table 2: Evaluation of models pretrained by 300k steps with AdamW and GO optimizers, then fine-tuned with AdamW optimizer. The GO optimizer effectively prevents forgetting during supervised fine-tuning while maintaining comparable performance during pretraining.
Model
OWT
NLL ($\downarrow$)
Lambada
PPL ($\downarrow$)
Lambada
ACC
HellaSwag SST2 WinoGrande ARC-E ARC-C
GPT-2-Official 3.11 38.6 30.6 38.9 56.7 51.3 43.1 17.2
AdamW-300k 2.94 38.3 32.4 40.2 51.4 50.9 44.1 18.2
GO-300k 2.97 40.8 32.9 38.0 57.7 51.6 41.2 18.9
AdamW-SFT 3.49 94.1 26.8 39.8 51.7 49.8 44.0 23.6
GO-SFT 3.37 80.1 28.3 37.8 57.1 51.1 42.8 21.8
To evaluate the effectiveness of the GO optimizer, we conducted pre-training and fine-tuning experiments on a GPT2-127M model, as detailed in Table˜2. The key findings are summarized below.
Low Computational Overhead of GO Optimizer in Pre-training. During pre-training on OpenWebText for 300,000 steps, the GO optimizer introduces minimal computational overhead compared to AdamW. As shown in Fig.˜5(a), the GO optimizer is initially slower due to its optimization over the entire parameter neighborhood, resulting in a negative log-likelihood (NLL) of 2.97 compared to AdamW’s 2.94, a marginal gap of 0.03. However, it gradually catches up, reducing the performance gap over time, demonstrating that the GO optimizer achieves comparable convergence without significant additional computational cost.
Implicit Biases Beyond Basin Enlargement. The GO optimizer exhibits implicit biases that extend beyond enlarging the most-case basin, leading to superior performance on specific benchmarks. As shown in Table˜2, the GO optimizer outperforms AdamW on LAMBADA accuracy (32.9 vs. 32.4), SST-2 (57.7 vs. 51.4), WinoGrande (51.6 vs. 50.9), and ARC-C (18.9 vs. 18.2) after 300,000 pre-training steps. These improvements suggest that the GO optimizer introduces beneficial inductive biases, enhancing model generalization on certain tasks.
Effective Prevention of Forgetting During Fine-tuning. In the fine-tuning phase on the Alpaca dataset using the AdamW optimizer, the GO-pre-trained model significantly mitigates forgetting of prior capabilities compared to the AdamW-pre-trained model. Although AdamW initially achieves better performance on some benchmarks during pre-training (e.g., OpenWebText NLL of 2.94 vs. 2.97), fine-tuning reveals substantial forgetting in the AdamW-pre-trained model, with OpenWebText NLL degrading to 3.49, compared to 3.37 for the GO-pre-trained model—a gap of 0.12. Similar trends are observed in LAMBADA perplexity (94.1 vs. 80.1) and accuracy (26.8 vs. 28.3), as shown in Table˜2, confirming that the GO optimizer effectively preserves prior capabilities during supervised fine-tuning.
D.6 About Norm Constraints in Landscape Visualization
@@
D.7 Norm of Common Fine-tuning
@@
We also test the $\ell_{2}$ norm of common post-training techniques to determine whether they are located within basins or guaranteed regions. As shown in Table˜3, we draw the following conclusions:
@@
Appendix E Proof of Eq.˜4 and Eq.˜5
E.1 Proof of Eq.˜5
In this section, we present a simplified proof of Eq.˜5 by adapting Lemma 2 from Salman et al. (2019). We first restate this lemma:
Lemma E.1
For any function $f:\mathbb{R}^{d}\to\mathbb{R}$, the map $\bm{x}\to\Phi^{-1}(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[f(\bm{x}+\bm{\epsilon})])$ is at most 1-Lipschitz.
By setting $J_{\mathcal{D}}$ as $f$ and $\bm{\theta}$ as $\bm{x}$, we establish that, for any benchmark function $J_{\mathcal{D}}:\mathbb{R}^{d}\to\mathbb{R}$, the map $\bm{\theta}\to\Phi^{-1}(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})])$ is 1-Lipschitz. Consequently, we consider the function:
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)$ $=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\sigma\cdot\bm{\epsilon})]\right)$
$=\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}\left(\sigma\cdot(\frac{\bm{\theta}}{\sigma}+\bm{\epsilon})\right)]\right),$
which is at most 1-Lipschitz with respect to $\frac{\bm{\theta}}{\sigma}$, implying that it is at most $\frac{1}{\sigma}$-Lipschitz with respect to $\bm{\theta}$. Thus, we obtain:
$\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\right)\geq\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}.$
By applying $\Phi$ to both sides, we derive:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$
which matches the result of Eq.˜5.
E.2 Proof of Eq.˜4
Researchers have provided various methods to prove Eq.˜4, e.g., through the Neyman-Pearson Lemma (Cohen et al., 2019), knapsack algorithm (Chen et al., 2025a), and Lipschitz properties (Salman et al., 2019; Chen et al., 2024). In this section, we deduce Eq.˜4 from Eq.˜5, avoiding repetition of previous proofs and providing a more direct connection between these theorems.
$\quad\bm{\theta}\to\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\text{ is $\frac{1}{\sigma}$-Lipschitz}$
$\iff\left\|\nabla_{\bm{\theta}}\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right\|\leq\frac{1}{\sigma}$
$\iff\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\Phi^{\prime}\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)\right)}\leq\frac{1}{\sigma}$
$\implies\frac{\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|}{\frac{1}{\sqrt{2\pi}}}\leq\frac{1}{\sigma}$
$\implies\left\|\nabla_{\bm{\theta}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right\|\leq\frac{1}{\sqrt{2\pi}\sigma}$
$\implies\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}[J_{\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\text{ is at most }\frac{1}{\sqrt{2\pi}\sigma}\text{-Lipschitz}$
E.3 Comparison of Eq.˜4 and Eq.˜5
@@
Beyond the fact that Eq.˜5 is strictly stronger than Eq.˜4 (i.e., the Lipschitz constant of Eq.˜5 at each point is strictly less than $\frac{1}{\sqrt{2\pi}\sigma}$), there is also a significant difference. Eq.˜5 indicates that the higher the clean accuracy $p_{A}$, the smaller the Lipschitz constant, and thus the greater the robustness; when $p_{A}=1$, the Lipschitz constant becomes zero, making forgetting extremely unlikely. Conversely, the lower the clean accuracy $p_{A}$, the larger the Lipschitz constant, and the more likely performance degradation occurs. This eliminates the forgetting-learning tradeoff even along the worst-case direction, i.e., as long as the performance on old tasks exceeds that on new tasks, the smoothness constraints make it more likely to learn new tasks than to forget old tasks, and the performance degradation on old tasks is always less than the maximum performance gain on new tasks.
Appendix F Limitations and Future Work
In this work, we discovered the emergence of basins: as models grow larger, they exhibit absolute robustness to Gaussian noise within a certain range, resulting in the formation of basins (Sec.˜3.2). Based on this, we conjecture that as long as benign fine-tuning occurs within the basin, the model will not forget existing capabilities. The reasoning for this conjecture is: If random directions are minimally harmful within a certain range, forming a basin of stability, benign fine-tuning directions should be no more harmful than random perturbations. We experimentally validated that benign fine-tuning behaves this way (Sec.˜3.4). More rigorously, we used randomized smoothing techniques to prove that, if the model has a $\sigma$-basin, performance degradation during benign fine-tuning can be bounded, providing a rigorous theoretical explanation for the conjecture (Sec.˜4.1). Additional contributions include: analyzing how parameter space robustness provides loose guarantees against jailbreaking attacks (Sec.˜4.2), exploring whether a regularization term could constrain supervised fine-tuning (SFT) within the basin in the future (Sec.˜4.3), and conducting a preliminary investigation into simple methods to enlarge the basin (Sec.˜5).
F.1 Randomized Smoothing Regime
As models grow larger, we draw the following conclusions:
  • As models grow larger, their basins naturally expand, increasing the theoretically guaranteed radius (which is $O(\sigma)$) for subsequent fine-tuning, making it harder to compromise prior capabilities.
  • As models and their basins grow larger, the regions within the basin exhibit greater expressive power, facilitating the acquisition of new capabilities.
We define the region where the model is guaranteed to preserve prior capabilities while learning new capabilities as the Randomized Smoothing Regime (RS Regime). The region where the model is guaranteed to preserve prior capabilities has a size of $O(\sigma)$ (see Sec.˜4.1). Taking Rademacher complexity as a metric for expressive power, for the expressive power to exceed $k$, we require $O(d\sigma^{2})\geq k\iff\sigma\geq O\left(\sqrt{\frac{k}{d}}\right)$. Thus, the RS Regime is $0,O(\sigma)]\cap\left[O\left(\sqrt{\frac{k}{d}}\right),+\infty\right)$. As models grow larger, $O(\sigma)$ continues to grow, while $O\left(\sqrt{\frac{k}{d}}\right)$ continues to decrease. Consequently, the RS Regime, representing the theoretically guaranteed region for preserving prior capabilities and acquiring new capabilities, continuously expands. When models have infinitely many parameters, all fine-tuning may lie within the RS Regime, analogous to the Neural Tangent Kernel (NTK) regime (Jacot et al., 2018), where infinitely wide models behave like kernel methods. Future work could explore how quickly models enter the RS Regime as their size increases and identify when this transition occurs.
F.2 Larger Scale Pre-training Studies
Although we conducted pre-training studies using the GO optimizer, the scale is limited to the NanoGPT level, and several questions remain unanswered:
  • As the model size increases, does it become easier or harder to expand the basin? Will the maximum basin size increase or decrease? Will the optimization speed increase or decrease? Addressing these questions may require a scaling law for basin size with respect to the number of training data and parameters.
  • As the data composition becomes more complex and the number of parameters increases, does the GO optimizer still introduce implicit biases beyond enlarging basins, as observed in Sec.˜D.5? Will these biases remain benign or become harmful?
In the future, we plan to conduct pre-training studies on the GO optimizer at a much larger scale, e.g., on 7B models and 10T training data.